OpenStudy (anonymous):

How to find value of discriminant & number of solutions?? 5x² - 4x + 6 = 0

OpenStudy (paki):

use formula of discriminants... and solve... b^2-4ac.... here a=5...b=-4...c=6

OpenStudy (anonymous):

What is the formula of discriminants?

OpenStudy (paki):

b^2-4ac

OpenStudy (anonymous):

Ok So it's 16=120 which would be 7.5 ?

OpenStudy (wiggler):

If the discriminant is more than zero, the equation has two solutions, if it is equal to zero there is one solution, and if it's less than zero there are no solutions.

OpenStudy (paki):

how 120... please show your work....

OpenStudy (the_fizicx99):

If it's equal to zero there will be two solutions, identical to each other.

OpenStudy (wiggler):

Since your discriminant is 7.5 (more than zero) there are two solutions.

OpenStudy (anonymous):

@paki I did b² (16) = 4 x 5 (a) x 6 (c) which was 120

OpenStudy (paki):

b^2-4ac= (-4)^2-4(5)*(6)

OpenStudy (paki):

when solve it will give us -104

OpenStudy (the_fizicx99):

\(\ \sf \sqrt{b^2 - 4ac} \) \(\ \sf A =\color{green}{ 5}\) \(\ \sf B =\color{blue}{-4}\) \(\ \sf C =\color{green}{ 6}\) \(\ \sqrt{ (\color{blue}{-4}) -4(\color{green}{5)(6)}} \) \(\ \sf \sqrt{16 -4(30)}\) \(\ \sf \sqrt{16 - 120}\) Now you get it? Continue :D Remember the rules for discriminates; If the discriminant is more than zero, the equation has two solutions, if it is equal to zero there is two solutions, each solution is identical, if it's less than zero there are no solutions.

OpenStudy (anonymous):

Oh sorry I dint see the minus sign! 16 -120 = -104

OpenStudy (anonymous):

And it's no solution

OpenStudy (the_fizicx99):

The reason there are no solutions is because the vertex doesn't touch the x axis 

OpenStudy (paki):

@Bloom23 so what will be the answer now :_)

OpenStudy (paki):

yup, no solution... good...

OpenStudy (the_fizicx99):

There is no solutions because you get a negative number, no number times itself is negative, 4*4 = 16, -4*-4 = 16, -4 * 4 isn't times itself. You would get a complex solutions i√104

OpenStudy (anonymous):

Ok thank you! @paki @tHe_FiZiCx99 @wiggler

OpenStudy (paki):

pleasure :)

OpenStudy (the_fizicx99):

You're welcome :)